Circular No. 8267

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 A1 B. A. Skiff, Lowell Observatory, reports the discovery of a comet found

on CCD images taken by himself (discovery observation tabulated below) in the course of the LONEOS program with the 0.59-m Schmidt telescope. The LONEOS images showed stellar appearance, but the motion led Skiff to request that a pair of 5-min R-band CCD exposures be taken of the object by H. R. Miller at the Perkins 1.8-m reflector, and these images show a 3′′ well-condensed coma and a 12′′ tail in p.a. 290^{o}. To get additional observations, the object’s ephemeris was then posted on the NEO Confirmation Page, which resulted in CCD images taken by J. Young (Table Mountain, CA, 0.6-m reflector) on Jan. 13.5 UT that show a 4′′ coma with a very faint 10′′ tail in p.a. 262^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Jan. 13.27307  8^{h}45^{m}26.^{s}75  +18^{o}11′50′′.2  18.4

Additional astrometry (including prediscovery observations on 2003 Dec. 18 from Spacewatch at Kitt Peak and on 2004 Jan. 5 from the Catalina Sky Survey), the following orbital elements, and an ephemeris appear on MPEC 2004-A51.

T = 2005 Mar. 4.815 TT  ω = 38.^{o}037
e = 0.32868  Ω = 125.304 }2000.0
q = 5.31259 AU  i = 10.920
a = 7.91368 AU n^{o} = 0.044273 P = 22.26 years

NOVA IN M32 J. D. Neill, American Museum of Natural History and Columbia Uni-

versity, reports the discovery of an apparent nova on V -band CCD images taken on Jan. 5.120 (mag 17.2), 7.120 (mag 17.3), and 9.121 UT (mag 17.4) with a 0.81-m reflector at Tenagra Observatory (near Nogales, AZ) in the course of a nova patrol of M31 companions. The new object is located at α = 0^{h}42^{m}44^{s}.991, δ = +40^{o}53′04′′.76 (equinox 2000.0), which is ≈ 35′′.7 east and 69′′.8 north of the center of M32. The new object is not present on Neill’s images from 2003 Dec. 29 (limiting mag 20.0), 28, and 27 or on Palomar Sky Survey red and blue plates. Spectra taken with the 2.4-m Hiltner Telescope (+ Modspec) at MDM Observatory by J. Thorstensen on Jan. 12.101 show strong, broad Balmer lines and Fe II lines, confirming the object as a slow nova (consistent with the observed fading rate).

2004 January 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green